# Symmetry-enforced Moiré Topology

### v1

@@ Abstract
Topological flat bands in two-dimensional (2D) moiré materials have emerged as promising platforms for exploring the interplay between topology and correlation effects. However, realistic calculations of moiré band topology using density functional theory (DFT) are computationally inefficient due to the large number of atoms in a single moiré unit cell. In this work, we propose a systematic scheme to predict the topology of moiré bands from atomic symmetry data and moiré symmetry group, both of which can be efficiently extracted from DFT. Specifically, for $\Gamma$-valley electron gases, we find that certain combinations of atomic symmetry data and moiré symmetry groups can enforce nontrivial band topology in the low-energy moiré bands, as long as the moiré band gap is smaller than the atomic band splitting at the moiré Brillouin zone boundary. This symmetry-enforced nontrivial moiré topology, including both topological insulators and topological semimetals, is robust against various material-specific details such as the precise form and strength of the moiré potential or the exact twist angle. By exhaustively scanning all 2D atomic symmetry data and moiré symmetry groups, we identify 197 combinations that can yield symmetry-enforced nontrivial moiré topology, and we verify one such combination using a moiré model with cubic Rashba spin-orbit coupling. Our approach is generalizable to other valleys and provides a useful guideline for experimental efforts to discover and design new topologically nontrivial moiré materials.
@@ I Introduction
Moiré materials [1, 2] have emerged as one of the most important platforms for exploring exotic strongly correlated physics, primarily because they host energy bands that are (nearly) flat and topological. The most recent prominent example is twisted bilayer MoTe2, which has been extensively studied both experimentally [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17] and theoretically [18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58], where the flat bands with nonzero Chern numbers give rise to fractional Chern insulators [59, 60, 61, 62, 63] at fractional fillings. The small bandwidth of the moiré bands in moiré materials naturally arises from the large moiré unit cell, which folds the non-moiré bands (i.e., atomic bands) hundreds of times, leading to energy bands that have bandwidths much smaller than their non-moiré counterparts. However, the nontrivial topology of the flat bands is not necessarily guaranteed. There is one known example of enforcing the nontrivial topology of the low-energy moiré bands based on the non-moiré band topology—twisted bilayer graphene (and related graphene-based systems) [1, 64, 2, 65, 66, 67, 68, 69, 70]. Yet, this case is quite special: one needs both the $C_{2}\mathcal{T}$ symmetry (i.e., combination of two-fold rotation $C_{2}$ and time reversal (TR) symmetry $\mathcal{T}$) and an effective normal-state particle-hole symmetry [69], which, especially the latter, are not common beyond graphene-based systems. Indeed, the nontrivial topology of the flat bands in twisted bilayer MoTe2 is not guaranteed—the Chern numbers of the bands are sensitive to material parameters such as twisted angles [42, 47, 55, 39].
@@
In this work, we successfully establish such a principle. Specifically, for $\Gamma$-valley electron gas subject to weak moiré potential, we find that certain combinations of atomic symmetry group, atomic symmetry representation at $\Gamma$ and moiré symmetry group can enforce nontrivial topology of the low-energy moiré bands, where nontrivial topology includes both topological insulator and topological semi-metal phases. Here the low-energy moiré bands refer to the moiré bands that are closest to the charge neutrality, i.e., conduction bottom bands or valence top bands. We refer to such enforced nontrivial topology as symmetry-enforced moiré topology. Weak moiré potential means that the strength of the moiré potential is weaker than the atomic band splitting, which is a realistic condition at $\Gamma$ valley satisfied by known examples such as twisted bilayer MoTe2, as discussed in Sec.V. In other words, for those combinations of atomic symmetry data and moiré symmetry group, the phase diagram of the low-energy bands of the moiré model only includes topological insulator and semi-metal phases—no topologically trivial phases. By systematically analyzing all possible combinations, we identify 197 combinations that can lead to symmetry-enforced nontrivial moiré topology, with 16 of them for the topological insulating phase in the phase diagram. The general principle is explicitly verified in a moiré model with cubic Rashba spin-orbit coupling (SOC). Clearly, our principle does not rely on precisely-tuned values of moiré parameters such as twist angle and moiré potential strength. It is also independent of the form of moiré potential, in contrast to Refs.[71, 72].
We note that the moiré topological semi-metal phases may also be a fertile ground for exotic strongly-correlated phases. It is because electron-electron interactions may open a gap between topological metallic bands and drive the system into topologically ordered states. A closely related example is the emergence of fractional Chern insulators in multilayer rhombohedral graphene/hexagonal boron nitride (hBN) heterostructures at various electron fillings [73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83], where the moiré bands are nearly gapless at the single-particle level [84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 71, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104].
@@ II Illustration of Principle
Now we resume the nonzero moiré strength $\Delta_{M}$. We consider the weak moiré region where the energy splitting induced by the moiré potential is smaller than the atomic band splitting ($\Delta_{A}$) at the mBZ boundary. The moiré potential will leave the $\bar{\Gamma}_{7}$ irrep at $\Gamma_{M}$ and the $\bar{\text{M}}_{5}$ irrep at $\text{M}_{M}$ intact, as protected by symmetry. Nevertheless, the moiré potential will split the $\bar{\text{K}}_{4}$ and $\bar{\text{K}}_{6}$ bands at $\text{K}_{M}$. Focusing on the top two bands, there are two possibilities. One is that the top two bands carry the 2D $\bar{\text{K}}_{6}$ irrep, and the third band possesses the $\bar{\text{K}}_{4}$ irrep, resulting in the top two bands being isolated (Fig. 2(c)). In this case, the top two bands have nontrivial topology as the irreps $\bar{\Gamma}_{7}$, $\bar{\text{M}}_{5}$ and $\bar{\text{K}}_{6}$ can never appear simultaneously for an isolated set of trivial bands according to topological quantum chemistry [105, 106, 107] and symmetry indicator theory [108, 109, 110, 111, 112, 113]. The other possibility is that the top two bands are from one 1D $\bar{\text{K}}_{4}$ irrep and one component of the 2D $\bar{\text{K}}_{6}$ irrep. In this case, the top second band definitely touches the third band as they together form the $\bar{\text{K}}_{6}$ irrep at $\text{K}_{M}$. Therefore, regardless of which possibility, the low-energy moiré bands are either topologically insulating or topologically semi-metallic. Such nontrivial moiré topology is enforced by the atomic symmetry data and moiré symmetry group, as long as the moiré potential is weak compared to atomic band splitting.
@@ III Moiré cubic Rashba model
To illustrate our principle, we focus on the weak moiré region, where the moiré gap is smaller than the band splitting caused by Rashba SOC. By choosing reasonable parameter values (listed in supplementary material (SM) Sec.C1 [114].), we obtain the phase diagram as a function of $\Delta_{1}$ and $\Delta_{2}$ in Fig. 3(a). The color represents the direct band gap $E_{\Delta}$ between the second and third topmost moiré bands (see Fig. 3(b)) and the red dashed line separates the insulating region from the semi-metallic region for the top two moiré bands. (See more details in SM Sec.C1 [114].) We find that regardless of the value of the moiré parameter values, there are only two phases for the low-energy physics: (i) topological insulator phase where the top two bands are isolated with band irreps ($\bar{\mathbf{\Gamma}}_{7}$, $\bar{\textbf{M}}_{5}$, $\bar{\textbf{K}}_{6}$) and have $\mathds{Z}_{2}=1$ in Fig. 3(b), and (ii) topological semi-metal phase where the top two bands are enforced to be connected to other bands in Fig. 3(c) as the irrep $\bar{\textbf{K}}_{4}$ for the topmost band is 1D, while $\bar{\textbf{M}}_{5}$ and $\bar{\mathbf{\Gamma}}_{7}$ irreps are 2D. This is consistent with the principle that we discussed in Sec. II.
@@ IV Symmetry-Enforced Moiré Topology in 2D Plane Group
As discussed in Sec. II, atomic symmetry data includes the atomic symmetry group $G^{A}$, and its irrep $\Lambda^{A}_{\Gamma}$ furnished by the considered states at $\Gamma$, and we use $G$ to label the moiré symmetry group. Here $G^{A}$ is simply the point group at $\Gamma$, $G^{A}_{0}$, if there is no TR symmetry, but also includes the TR symmetry, denoted as $\mathcal{T}$, if it appears. Similarly, $G$ is simply the moiré plane group, $G_{0}$, if there is no TR symmetry, but also includes the TR symmetry if it appears. We focus on the case that the moiré potential does not cause extra TR breaking. For the specific case in Sec. II, we have $G^{A}=G^{A}_{0}\rtimes\mathcal{T}$ with $G^{A}_{0}=C_{6v}$, the irrep $\Lambda^{A}_{\Gamma}=\bar{\Gamma}_{7}$, and $G=G_{0}\rtimes\mathcal{T}$ with $G_{0}=P6mm$. Although we have the atomic symmetry group to be included by the moiré symmetry group in Sec. II, we do not assume such a relation in the general study in this section. The detailed steps for this general search are illustrated in SM Sec.A [114].
We went through all possible combinations of $G^{A}$, the irrep $\Lambda^{A}_{\Gamma}$, $G$, and the presence/absence of SOC, as elaborated in SM Sec.B [114]. The results are summarized in Tab. 1 for the spinful case with TR and in Tab. 2 for the spinless case without TR, where spinful (spinless) simply refers to the case with (without) SOC as indicated by the irrep notation with (without) a bar [115, 116, 117]. (The results of spinless cases with TR and spinful cases without TR are summarized in Tab. SI(c) and SIV(b) in the SM Sec.B [114] respectively. ) We find 58 combinations that lead to symmetry-enforced moiré topology in the spinful case, and 139 in the spinless case. Out of the 58 spinful combinations (139 spinless combinations), we find that there are 11 (5) combinations that allow for isolated topological bands at low-energy, while all other combinations only have topological semi-metal phases in the weak moiré potential limit. The symmetry-enforced gapped moiré topology has the following scenarios: (i) for the TR-preserving case, isolated topological moiré bands may carry nonzero $\mathds{Z}_{2}$ index in addition to other possibilities such as fragile topology; (ii) for the TR-breaking case, the topological moiré bands always have nonzero Chern number.
@@
On the other hand, symmetry-enforced moiré topological semimetals are much easier to be realized than symmetry-enforced moiré topological insulators. There are no particular constraints on the lattice type, or the simultaneous presence/absence of TR symmetry and SOC. Nevertheless, doublets at $\Gamma$ are still required to realize symmetry-enforced moiré topological semimetal, unless the moiré plane group is non-symmorphic [118, 119, 120, 121, 122].
V Conclusion and Discussion
We have proposed the concept of symmetry-enforced moiré topology, and list all combinations of atomic symmetry data and moiré symmetry group to realize it for $\Gamma$-valley electron gas subject to a weak moiré potential. We note that although the SOC term may have a high power of momentum, such as the $k^{3}$ term in Eq. 3, the weak moiré condition is still realistic. To show its realistic nature in our consideration, we consider the bands of twisted bilayer MoTe2 near $\Gamma$ valley. In monolayer MoTe2, the low-energy $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$ doublet at $\Gamma$ (included in Tab. 1) also has a $k^{3}$ band splitting term to the leading order. Yet, its atomic band splitting is around 7-11 meV at the edge of the mBZ of twisted bilayer MoTe2 with twist angle from 3.5 to 4, whereas the moiré band gap is about 3 meV [40], naturally satisfying the weak moiré condition. (See SM Sec.D [114] for details.) Moreover, the TR-breaking cases with the spinless models in Sec.C2 of SM [114] can also be applied to topological moiré magnon systems [123, 124, 125].

### v2

@@ Abstract
Topological flat bands in two-dimensional (2D) moiré materials have emerged as promising platforms for exploring the interplay between topology and correlation effects. However, realistic calculations of moiré band topology using density functional theory (DFT) are computationally inefficient due to the large number of atoms in a single moiré unit cell. In this work, we propose a systematic scheme to predict the topology of moiré bands from atomic symmetry data and moiré symmetry group, both of which can be efficiently extracted from DFT. Specifically, for $\Gamma$-valley electron gases, we find that certain combinations of atomic symmetry data and moiré symmetry groups can enforce nontrivial band topology in the low-energy moiré bands, as long as the moiré band gap is smaller than the atomic band splitting at the moiré Brillouin zone boundary. This symmetry-enforced nontrivial moiré topology, including both topological insulators and topological semimetals, is robust against various material-specific details such as the precise form and strength of the moiré potential or the exact twist angle. By exhaustively scanning all 2D atomic symmetry data and moiré symmetry groups, we identify 197 combinations that can yield symmetry-enforced nontrivial moiré topology, and we verify one such combination using a moiré model with cubic Rashba spin-orbit coupling. By screening the existing 2D material database, we currently identify 92 monolayer materials with (i) the low-energy bands near $\Gamma$ and (ii) the atomic symmetry data that belong to those combinations. Our approach is generalizable to other valleys and provides a useful guideline for experimental efforts to discover and design new topologically nontrivial moiré materials.
@@ I Introduction
Moiré materials [cao2018unconventional, bistritzer2011moire] have emerged as one of the most important platforms for exploring exotic strongly correlated physics, primarily because they host energy bands that are (nearly) flat and topological. The most recent prominent example is twisted bilayer MoTe2, which has been extensively studied both experimentally [cai2023signatures, zeng2023thermodynamic, park2023observation, xu2023observation, ji2024local, redekop2024direct, kang2024evidence, park2025ferromagnetism, an2025observation, jia2025anomalous, xu2025interplay, kang2025time, park2025observation, xu2025signatures, xia2025simulating] and theoretically [wu2019topological, yu2020giant, pan2020band, zhang2021electronic, li2021spontaneous, devakul2021magic, morales2023pressure, wang2024fractional, reddy2023fractional, qiu2023interaction, dong2023composite, wang2023diverse, goldman2023zero, morales2024magic, liu2024gate, xu2024maximally, reddy2023toward, song2024phase, wu2024time, yu2024fractional, abouelkomsan2024band, li2024electrically, jia2024moire, mao2024transfer, zhang2024polarization, wang2023topology, li2024contrasting, sheng2024quantum, reddy2024non, xu2025multiple, ahn2024non, wang2025higher, shen2024stabilizing, wang2024phase, kwan2024could, wu2024quantum, zaklama2025structure, zhang2024universal, qiu2025topological, gonccalves2025spinless, liu2025characterization], where the flat bands with nonzero Chern numbers give rise to fractional Chern insulators [neupert2011fractional, sheng2011fractional, regnault2011fractional, tang2011high, sun2011nearly] at fractional fillings. The small bandwidth of the moiré bands in moiré materials naturally arises from the large moiré unit cell, which folds the non-moiré bands (i.e., atomic bands) hundreds of times, leading to energy bands that have bandwidths much smaller than their non-moiré counterparts. However, the nontrivial topology of the flat bands is not necessarily guaranteed. There is one known example of enforcing the nontrivial topology of the low-energy moiré bands based on the non-moiré band topology—twisted bilayer graphene (and related graphene-based systems) [cao2018unconventional, cao2018correlated, bistritzer2011moire, song2019all, po2019faithful, ahn2019failure, tarnopolsky2019origin, song2021twisted, song2022magic]. Yet, this case is quite special: one needs both the $C_{2}\mathcal{T}$ symmetry (i.e., combination of two-fold rotation $C_{2}$ and time reversal (TR) symmetry $\mathcal{T}$) and an effective normal-state particle-hole symmetry [song2021twisted], which, especially the latter, are not common beyond graphene-based systems. Indeed, the nontrivial topology of the flat bands in twisted bilayer MoTe2 is not guaranteed—the Chern numbers of the bands are sensitive to material parameters such as twisted angles [zhang2024polarization, xu2025multiple, zhang2024universal, li2024electrically].
@@
In this work, we successfully establish such a principle. Specifically, for $\Gamma$-valley electron gas subject to weak moiré potential, we find that certain combinations of atomic symmetry group, atomic symmetry representation at $\Gamma$ and moiré symmetry group can enforce nontrivial topology of the low-energy moiré bands, where nontrivial topology includes both topological insulator and topological semi-metal phases. Here the low-energy moiré bands refer to the moiré bands that are closest to the charge neutrality, i.e., conduction bottom bands or valence top bands. We refer to such enforced nontrivial topology as symmetry-enforced moiré topology. Weak moiré potential means that the strength of the moiré potential is weaker than the atomic band splitting, which is a realistic condition at $\Gamma$ valley satisfied by known examples such as twisted bilayer MoTe2, as discussed in Sec.VI. In other words, for those combinations of atomic symmetry data and moiré symmetry group, the phase diagram of the low-energy bands of the moiré model only includes topological insulator and semi-metal phases—no topologically trivial phases. By systematically analyzing all possible combinations, we identify 197 combinations that can lead to symmetry-enforced nontrivial moiré topology, with 16 of them for the topological insulating phase in the phase diagram. The general principle is explicitly verified in a moiré model with cubic Rashba spin-orbit coupling (SOC). We further screen the existing non-moiré 2D material databases [vergniory2019complete, haastrup2018computational, jiang20242d], focusing on monolayers with plane groups P31m, P3m1, and P3. We identify 92 monolayer materials that have (i) low-energy bands near $\Gamma$ and (ii) the atomic symmetry data that belong to those combinations, as discussed in Sec. V.
Clearly, our principle does not rely on precisely-tuned values of moiré parameters such as twist angle and moiré potential strength. It is also independent of the form of moiré potential, in contrast to Refs.[crepel2025efficient, Lhachemiarxiv]. We note that the moiré topological semi-metal phases may also be a fertile ground for exotic strongly-correlated phases. It is because electron-electron interactions may open a gap between topological metallic bands and drive the system into topologically ordered states. A closely related example is the emergence of fractional Chern insulators in multilayer rhombohedral graphene/hexagonal boron nitride (hBN) heterostructures at various electron fillings [lu2024fractional, xie2025tunable, waters2025chern, lu2025extended, aronson2025displacement, zhou2024layer, han2024engineering, xiang2025continuously, wang2025electrical, li2025tunable, xie2025unconventional], where the moiré bands are nearly gapless at the single-particle level [park2023topological, dong2024theory, zhou2024fractional, dong2024anomalous, herzog2024moire, guo2024fractional, kwan2025moire, soejima2024anomalous, dong2024stability, xie2024integer, kudo2024quantum, crepel2025efficient, huang2024self, yu2024moir, huang2025fractional, das2024thermal, wei2025edge, zhou2025new, zeng2025berry, bernevig2025berry, li2025multiband, uchida2025non].
@@ II Illustration of Principle
Now we resume the nonzero moiré strength $\Delta_{M}$. We consider the weak moiré region where the energy splitting induced by the moiré potential is smaller than the atomic band splitting ($\Delta_{A}$) at the mBZ boundary. The moiré potential will leave the $\bar{\Gamma}_{7}$ irrep at $\Gamma_{M}$ and the $\bar{\text{M}}_{5}$ irrep at $\text{M}_{M}$ intact, as protected by symmetry. Nevertheless, the moiré potential will split the $\bar{\text{K}}_{4}$ and $\bar{\text{K}}_{6}$ bands at $\text{K}_{M}$. Focusing on the top two bands, there are two possibilities. One is that the top two bands carry the 2D $\bar{\text{K}}_{6}$ irrep, and the third band possesses the $\bar{\text{K}}_{4}$ irrep, resulting in the top two bands being isolated (Fig. 2(c)). In this case, the top two bands have nontrivial topology as the irreps $\bar{\Gamma}_{7}$, $\bar{\text{M}}_{5}$ and $\bar{\text{K}}_{6}$ can never appear simultaneously for an isolated set of trivial bands according to topological quantum chemistry [bradlyn2017topological, cano2021band, elcoro2021magnetic] and symmetry indicator theory [po2017symmetry, kruthoff2017topological, fu2007topological, fang2012bulk, po2020symmetry, lenggenhager2022universal]. The other possibility is that the top two bands are from one 1D $\bar{\text{K}}_{4}$ irrep and one component of the 2D $\bar{\text{K}}_{6}$ irrep. In this case, the top second band definitely touches the third band as they together form the $\bar{\text{K}}_{6}$ irrep at $\text{K}_{M}$. Therefore, regardless of which possibility, the low-energy moiré bands are either topologically insulating or topologically semi-metallic. Such nontrivial moiré topology is enforced by the atomic symmetry data and moiré symmetry group, as long as the moiré potential is weak compared to atomic band splitting.
@@ III Moiré cubic Rashba model
To illustrate our principle, we focus on the weak moiré region, where the moiré gap is smaller than the band splitting caused by Rashba SOC. By choosing reasonable parameter values (listed in supplementary material (SM) Sec.C1 [SM].), we obtain the phase diagram as a function of $\Delta_{1}$ and $\Delta_{2}$ in Fig. 3(a). The color represents the direct band gap $E_{\Delta}$ between the second and third topmost moiré bands (see Fig. 3(b)) and the red dashed line separates the insulating region from the semi-metallic region for the top two moiré bands. (See more details in SM Sec.C1 [SM].) We find that regardless of the value of the moiré parameter values, there are only two phases for the low-energy physics: (i) topological insulator phase where the top two bands are isolated with band irreps ($\bar{\mathbf{\Gamma}}_{7}$, $\bar{\textbf{M}}_{5}$, $\bar{\textbf{K}}_{6}$) and have $\mathds{Z}_{2}=1$ in Fig. 3(b), and (ii) topological semi-metal phase where the top two bands are enforced to be connected to other bands in Fig. 3(c) as the irrep $\bar{\textbf{K}}_{4}$ for the topmost band is 1D, while $\bar{\textbf{M}}_{5}$ and $\bar{\mathbf{\Gamma}}_{7}$ irreps are 2D. This is consistent with the principle that we discussed in Sec. II.
@@ IV Symmetry-Enforced Moiré Topology in 2D Plane Group
As discussed in Sec. II, atomic symmetry data includes the atomic symmetry group $G^{A}$, and its irrep $\Lambda^{A}_{\Gamma}$ furnished by the considered states at $\Gamma$, and we use $G$ to label the moiré symmetry group. Here $G^{A}$ is simply the point group at $\Gamma$, $G^{A}_{0}$, if there is no TR symmetry, but also includes the TR symmetry, denoted as $\mathcal{T}$, if it appears. Similarly, $G$ is simply the moiré plane group, $G_{0}$, if there is no TR symmetry, but also includes the TR symmetry if it appears. We focus on the case that the moiré potential does not cause extra TR breaking. For the specific case in Sec. II, we have $G^{A}=G^{A}_{0}\rtimes\mathcal{T}$ with $G^{A}_{0}=C_{6v}$, the irrep $\Lambda^{A}_{\Gamma}=\bar{\Gamma}_{7}$, and $G=G_{0}\rtimes\mathcal{T}$ with $G_{0}=P6mm$. Although we have the atomic symmetry group to be included by the moiré symmetry group in Sec. II, we do not assume such a relation in the general study in this section. The detailed steps for this general search are illustrated in SM Sec.A [SM].
We went through all possible combinations of $G^{A}$, the irrep $\Lambda^{A}_{\Gamma}$, $G$, and the presence/absence of SOC, as elaborated in SM Sec.B [SM]. The results are summarized in Tab. 1 for the spinful case with TR and in Tab. 2 for the spinless case without TR, where spinful (spinless) simply refers to the case with (without) SOC as indicated by the irrep notation with (without) a bar [aroyo2011crystallography, aroyo2006bilbao, aroyo2006bilbao1]. (The results of spinless cases with TR and spinful cases without TR are summarized in Tab. SI(c) and SIV(b) in the SM Sec.B [SM] respectively. ) We find 58 combinations that lead to symmetry-enforced moiré topology in the spinful case, and 139 in the spinless case. Out of the 58 spinful combinations (139 spinless combinations), we find that there are 11 (5) combinations that allow for isolated topological bands at low-energy, while all other combinations only have topological semi-metal phases in the weak moiré potential limit. The symmetry-enforced gapped moiré topology has the following scenarios: (i) for the TR-preserving case, isolated topological moiré bands may carry nonzero $\mathds{Z}_{2}$ index in addition to other possibilities such as fragile topology; (ii) for the TR-breaking case, the topological moiré bands always have nonzero Chern number.
@@
On the other hand, symmetry-enforced moiré topological semimetals are much easier to be realized than symmetry-enforced moiré topological insulators. There are no particular constraints on the lattice type, or the simultaneous presence/absence of TR symmetry and SOC. Nevertheless, doublets at $\Gamma$ are still required to realize symmetry-enforced moiré topological semimetal, unless the moiré plane group is non-symmorphic [michel1999connectivity, michel2001elementary, watanabe2015filling, watanabe2016filling, vergniory2017graph].
V Material candidates
We now discuss the material candidates. As our framework only needs information of the monolayers, we can screen through the existing non-moiré 2D material databases [haastrup2018computational, vergniory2019complete, jiang20242d] for suitable monolayer 2D materials for our proposal.
We first pick out monolayer spin-orbit coupled materials in which (i) VBM or CBM is located at $\Gamma$ and (ii) there are no symmetries that will force the atomic band splitting to be zero along at least one line in the BZ. Particularly, we focus on atomic plane groups P31m, P3m1 and P3, in which 92 monolayer materials are identified with their atomic symmetry data listed in Tab. 1. In particular, for 71 of them, the required moiré symmetry group is P3 according to Tab. 1, which can be naturally realized by a small angle twisted homobilayer structure.
We show two examples, BiBrTe and BiFTe, in Fig. 4. For these two examples, the atomic symmetry group at $\Gamma$ is $G_{0}^{A}=C_{3v}$ and the irreps of VBM are $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$, which belong to Tab. 1. The required moiré symmetry group P3 naturally arise in the twist homobilayer of either of the two examples (See more discussion in Sec.D in SM [SM].
Figure 4: (a) Band dispersion of BiBrTe. (b) Band dispersion of BiFTe. The irreps of VBM for both materials are $\bar{\Gamma}_{4}\bar{\Gamma}_{5}$. The inset figures show a zoom-in of the regions inside the dashed boxes.
VI Conclusion and Discussion
We have proposed the concept of symmetry-enforced moiré topology, and list all combinations of atomic symmetry data and moiré symmetry group to realize it for $\Gamma$-valley electron gas subject to a weak moiré potential. 92 candidate monolayer materials are identified based on a selective search.